OpenStudy (anonymous):

One side of a triangle is 4cm longer than the shortest side and 2 cm shorter than the longest side. The perimeter is 38 cm. find the dimensions of the triangle.

OpenStudy (midhun.madhu1987):

Let the side be x. Then shortest side = x - 4 Longest side = x + 2 Perimeter = 38 Hence x + x - 4 + x + 2 = 38 Find x.

OpenStudy (anonymous):

Since you only know the proportions of each side, use x for now. Sounds like the side they're describing is the middle one in terms of length. So, x is the side they are describing... It is 4cm longer than the shortest side. So the shortest side equals x, only minus 4cm (x-4). It is also 2cm shorter than the longest. The same reasoning the longer side is (x+2). The perimeter is the sum of all three sides. x + (x-4) + (x+2) =38 x+x-4+x+2=38 3x-2=38 3x=40 x=40/3=13.3333 The other sides are x-4=9.3333 and x+2=15.3333 So, it'd be 9.3, 13.3, 15.3.